# Functional normalizing flow for statistical inverse problems of partial differential equations

### v1

@@ 3 Conditional Normalizing Flow in Function Space
While Section 2 establishes the theoretical framework for functional normalizing flow, the practical application of Algorithm 1 is hindered by the computational cost of repeatedly solving forward problems. To address this limitation, strategies for reducing the algorithm’s iteration are essential.
In order to mitigate the computational burden, we propose conditional functional normalizing flow, which involve pre-training a model to map measurement data and their corresponding measurement points to the parameters of a functional normalizing flow, enabling efficient estimation of posterior for new measurement.
Our objective is to learn the map from measurement information (measurement data $\bm{d}$ and their corresponding measurement points $\bm{x}$) to the parameters of a functional normalizing flow model, enabling approximation of posterior corresponding to the given measurement information. Following model training, the conditional functional normalizing flow can efficiently generate approximate posterior for new measurement information. While this provides an initial estimate, further refinement can be achieved through fine-tuning the model on specific measurement information, leveraging the initial estimate as a starting point. This approach significantly reduces the number of iterations required compared with training from a random initialization.
3.1 Basic Setting of Conditional Normalizing Flow
Conditional functional normalizing flow aim to solve a class of inverse problems using a unified model. In constructing such a model, the position of the measurement points $\bm{x}$ and the measurement data $\bm{d}$ play an important role. When employing a conditional network for solving inverse problems, both $\bm{x}$ and $\bm{d}$ are considered variable quantities. Therefore, we need to reformulate the framework proposed in Section 2. Let $\bm{x}=(x^{1},x^{2},\ldots,x^{N_{\bm{d}}})$ represents the measurement points of the solution, and let $\bm{d}=(w(x^{1}),w(x^{2}),\ldots,w(x^{N_{\bm{d}}}))$ denotes the measurement data, where $w$ is the function $\mathcal{G}(u)$.
Let $\mathcal{G}:\mathcal{H}_{u}\rightarrow\mathcal{H}_{w}$ represents the solution operator of the corresponding PDE, where $\mathcal{H}_{u}$ is the parameter space, and $\mathcal{H}_{w}$ is the solution space. The measurement locations, denoted by $\bm{x}$, influence the observation operator $\mathcal{S}_{\bm{x}}$, formally defined as:
$\mathcal{S}_{\bm{x}}:\mathcal{H}_{w}\rightarrow\mathbb{R}^{N_{\bm{d}}},$
@@ 3.1 Basic Setting of Conditional Normalizing Flow
Then, the problem can be formulated as
$\bm{d}=\mathcal{S}_{\bm{x}}\mathcal{G}(u)+\epsilon,$
where $\bm{d}\in\mathbb{R}^{N_{\bm{d}}}$ is the measurement data, $u\in\mathcal{H}_{u}$ is the interested parameter, and $\bm{\epsilon}$ is a Gaussian random vector with zero mean and covariance matrix $\bm{\Gamma}_{\text{noise}}:=\tau^{-1}\textbf{I}$ ($\tau$ is a fixed positive number, I denotes the $N_{\bm{d}}$-dimensional identity matrix), which means
@@
Our objective is to infer the possible values of the function parameter $u$ based on measurement data $\bm{d}$ and their corresponding measurement points $\bm{x}$. After incorporating the measurement points $\bm{x}$, the Bayes’ formula is given by
@@
where $\mu_{(\bm{x},\bm{d})}$ is the posterior, the potential function $\Phi_{(\bm{x},\bm{d})}:\mathcal{H}_{u}\rightarrow\mathbb{R}$ is defined as
$\Phi_{(\bm{x},\bm{d})}(u)=\frac{1}{2}\lVert\bm{d}-\mathcal{S}_{\bm{x}}\mathcal{G}(u)\rVert^{2}_{\bm{\Gamma}_{\text{noise}}},$
with $\lVert\cdot\rVert_{\bm{\Gamma}_{\text{noise}}}:=\lVert\bm{\Gamma}_{\text{noise}}^{-1/2}\cdot\rVert$, and $Z_{\mu}^{(\bm{x},\bm{d})}$ is the normalization constant given by
@@
The goal of conditional functional normalizing flow is to approximate the posterior of Bayes’ formulas corresponding to different measurement information $(\bm{x},\bm{d})$, but with the same forward operator $\mathcal{G}$. We hope that, after entering the measurement information $(\bm{x},\bm{d})$ into the neural network, we can directly obtain an approximate measure of the posterior $\mu_{(\bm{x},\bm{d})}$ of the corresponding Bayes’ formula (3.1).
It is crucial to note that the measurement information $(\bm{x},\bm{d})$ is dynamic and heterogeneous. Variations in the number, positions, and values of measurement points lead to varying dimensions and scales of the information. To enable the application of neural networks, we aim to transform this diverse data into a fixed-length vector.
@@
$\mathcal{S}_{\bm{x}}^{*}:\mathbb{R}^{N_{\bm{d}}}\rightarrow\mathcal{H}_{w}^{*}.$
@@
Unlike functional normalizing flow introduced in Section 2, we propose a novel approach where the model parameters for each measurement information $(\bm{x},\bm{d})$ are generated directly by a neural network $\mathcal{N}_{\lambda}(\bm{v})$, rather than being learned through training. In other words, once the parameters $\lambda$ of the neural network $\mathcal{N}_{\lambda}(\bm{v})$ are trained, we can directly generate the parameters of the corresponding flow model for any given measurement information $(\bm{x},\bm{d})$. Next, we will delve into the details of training the parameters $\lambda$ of the neural network $\mathcal{N}_{\lambda}(\bm{v})$.
Given a dataset $D_{train}=\{(\bm{x}_{1},\bm{d}_{1}),(\bm{x}_{2},\bm{d}_{2}),\ldots,(\bm{x}_{N_{train}},\bm{d}_{N_{train}})\}$, where $({\bm{x}}_{i},{\bm{d}}_{i})$ represents a pair of measurement data $\bm{d}_{i}$ and its corresponding measurement points $\bm{x}_{i}$ for $i=1,2,\ldots,N_{train}$, the neural network $\mathcal{N}_{\lambda}(\bm{v})$ can be trained. For a specific measurement information pair $(\bm{x}^{*},\bm{d}^{*})$, the parameters of the corresponding functional normalizing flow are given by $\theta^{*}=\mathcal{N}_{\lambda}(\bm{v}^{*})$, where $\bm{v}^{*}=\mathcal{W}(\bm{x}^{*},\bm{d}^{*})$. As shown in (3.1), we denote the posterior corresponding to $(\bm{x}^{*},\bm{d}^{*})$ as $\mu_{(\bm{x}^{*},\bm{d}^{*})}$. We hypothesize that the flow model $f_{\theta^{*}}$ can transform the prior $\mu_{0}$ into a measure $\nu_{\lambda}(\bm{v}^{*})$ that closely approximates the posterior $\mu_{(\bm{x}^{*},\bm{d}^{*})}$. Consequently, we expect the Kullback-Leibler divergence $D_{\text{KL}}(\nu_{\lambda}(\bm{v}^{*})||\mu_{(\bm{x}^{*},\bm{d}^{*})})$ to be small.
Let $\{\bm{v}_{1},\bm{v}_{2},\ldots,\bm{v}_{N_{train}}\}$ be the corresponding vectors of the dataset $D_{train}$. We anticipate that for any $i=1,2,\ldots,N_{train}$, $D_{\text{KL}}(\nu_{\lambda}(\bm{v}_{i})||\mu_{(\bm{x}_{i},\bm{d}_{i})})$ will be small. The overall error function can be expressed as
@@
For a detailed description of the algorithm, please refer to Algorithm 2.
@@
1:  Initialize the neural network parameter $\lambda$ to $\lambda_{0}$, the learning rate schedule to $\alpha_{k}$, the training steps to $K$, the sampling size to $M$, and the training dataset to $D_{train}=\{(\bm{x}_{1},\bm{d}_{1}),(\bm{x}_{2},\bm{d}_{2}),\ldots,(\bm{x}_{N_{train}},\bm{d}_{N_{train}})\}$. Then, initialize the iteration counter $k$ to 00;
@@
3.2 Further Training for Specific Data
In Subsection 3.1, we train a neural network $\mathcal{N}_{\lambda}(\bm{v})$ of the conditional functional normalizing flow. For a given measurement information pair $(\bm{x}^{*},\bm{d}^{*})$, the corresponding vector $\bm{v}^{*}=\mathcal{W}(\bm{x}^{*},\bm{d}^{*})$ is input to this trained network, yielding parameters $\theta^{*}=\mathcal{N}_{\lambda}(\bm{v}^{*})$ for the functional normalizing flow. The flow model $f_{\theta^{*}}$ can then transform the prior $\mu_{0}$ into an approximate measure $\nu_{\lambda}(\bm{v}^{*})$ of the posterior $\mu_{(\bm{x}^{*},\bm{d}^{*})}$. However, due to the inherent limitations of the neural network $\mathcal{N}_{\lambda}(\bm{v})$, this initial estimate may deviate from the true posterior. Consequently, the primary objective of this subsection is to develop a method for refining the approximate posterior $\nu_{\lambda}(\bm{v}^{*})$ to align it more closely with the true posterior $\mu_{(\bm{x}^{*},\bm{d}^{*})}$. This refinement process is crucial for enhancing the accuracy of our inference.
To refine the approximate posterior $\nu_{\lambda}(\bm{v}^{*})$, we propose a retraining approach. For a specific measurement information pair $(\bm{x}^{*},\bm{d}^{*})$, the corresponding vector $\bm{v}^{*}=\mathcal{W}(\bm{x}^{*},\bm{d}^{*})$ serves as input to the trained neural network $\mathcal{N}_{\lambda}(\bm{v})$ to obtain initial parameters $\theta_{0}=\mathcal{N}_{\lambda}(\bm{v}^{*})$. Subsequently, the inverse problem corresponding to $(\bm{x}^{*},\bm{d}^{*})$ is retrained using Algorithm 1 with $\theta_{0}$ as a starting point. For a detailed overview of the entire process, refer to Algorithm 3. Compared to the random initialization, this approach significantly reduces the number of iterations required for training a functional normalizing flow, as demonstrated numerically in Section 5.
Algorithm 3 Retraining Functional Normalizing Flow
1:  For a new measurement information $(\bm{x}^{*},\bm{d}^{*})$, we change it into the corresponding vector $\bm{v}^{*}=\mathcal{W}(\bm{x}^{*},\bm{d}^{*})$, and put the vector into the trained neural network to get the parameters $\theta_{\lambda}=\mathcal{N}_{\lambda}(\bm{v}^{*})$;
2:  Let $\theta_{0}=\theta_{\lambda}$ be the initial value of $\theta$, set the sampling size $N$, the iterative number $K$, the learning rate schedule $\alpha_{k}$, and initialize the iteration counter $k=0$;
3:  repeat
4:     Sampling $N$ functions $\{u_{i}\}_{i=1}^{N}$ from the prior $\mu_{0}$;
5:     Update the parameters $\theta_{k+1}=\theta_{k}-\alpha_{k}\nabla_{\theta_{k}}L(\theta_{k})$ with
$\nabla_{\theta_{k}}L(\theta_{k})$ $=\nabla_{\theta_{k}}D_{\text{KL}}(\mu_{f_{\theta_{k}}}||\mu_{(\bm{x}^{*},\bm{d}^{*})})$
$\approx\frac{1}{N}\sum\limits_{i=1}^{N}\nabla_{\theta_{k}}\ln\left(\frac{d\mu_{f_{\theta_{k}}}}{d\mu_{0}}\left(f_{\theta_{k}}(u_{i})\right)\right)-\frac{1}{N}\sum\limits_{i=1}^{N}\nabla_{\theta_{k}}\ln\left(\frac{d\mu_{(\bm{x}^{*},\bm{d}^{*})}}{d\mu_{0}}\left(f_{\theta_{k}}(u_{i})\right)\right).$
6:     $k=k+1$;
7:  until $k=K$.
8:  Return the final result $\theta=\theta_{K}$.
@@ 4 Numerical Examples of Functional Normalizing Flow
To demonstrate the effectiveness of our novel framework for solving inverse problems, we present two illustrative examples, i.e., the simple elliptic equation and the steady-state Darcy flow equation. These examples enable a straightforward visualization of the approximate posterior distribution and facilitate direct comparisons with benchmark results obtained from the well-established preconditioned Crank-Nicolson (pCN) algorithm [17]. All of the codes for the functional normalizing flow are available at: https://github.com/jjx323/FunctionalNormalizingFlow.
In the first example, given the linear nature of the inverse problem, we utilize functional Householder flow and functional projected transformation flow as our flow models. For the second example, which involves a nonlinear inverse problem, we employ functional planar flow and functional Sylvester flow. These flow models are specifically designed to capture intricate nonlinear relationships, enabling the transformation of a Gaussian distribution into a non-Gaussian distribution that more closely approximates the posterior distribution associated with the nonlinear inverse problem.
@@ 4.1 Simple Elliptic Equation
$\begin{split}-\alpha\Delta w+w&=u,\quad\text{in}\ \Omega,\\ \frac{\partial w}{\partial\bm{n}}&=0,\quad\text{on}\ \partial\Omega,\end{split}$ (4.1)
where $\Omega=(0,1)\subset\mathbb{R}$, $\alpha>0$ is a positive constant, and $\bm{n}$ denotes the outward unit normal vector. The forward operator is defined as
@@
    Let the domain $\Omega$ be an interval $(0,1)$ with $\partial\Omega=\{0,1\}$, and the measurement data are assumed to be $\{w(x^{i})|i=1,2,\ldots,10\}$, where $x^{i}=\frac{i}{10}$.
@@
    In order to avoid inverse crime [27], the data is generated on a fine mesh with the number of grid points equal to $10^{4}$. And we use different sizes of mesh $n=\{50,75,100,200,300\}$ in the inverse stage.
4.1.1 Functional Householder Flow
In this subsection, we will show the effectiveness of functional Householder flow and the discretization invariance of this model. We assume that the data produced from the underlying ground truth
@@ 4.1.1 Functional Householder Flow
To demonstrate the superiority of our proposed algorithm, we conduct a comparative study with the pCN algorithm. It has been theoretically established that the pCN algorithm can generate samples from the true posterior distribution [17]. Consequently, we treat the samples obtained from pCN as benchmark results for comparison. However, as a Markov chain Monte Carlo (MCMC) method, pCN samples often exhibit high autocorrelation. To acquire a sufficient number of independent samples, numerous iterations are typically required, and the computational cost of pCN scales directly with the desired number of independent samples.
In contrast, the computational burden of NF-iVI primarily lies in the training of the flow model. Once trained, the generated samples are mutually independent, which enables efficient sampling. To quantify this advantage, we employ the effective sample size (ESS) as a metric. ESS is a statistical concept used to quantify the number of independent samples within a given dataset. For a more detailed explanation, refer to [43]. Assume we have a series of sample functions $u_{1},u_{2},\ldots,u_{n}$, and we are interested in the values of $f(u_{i})$, where $f$ is a functional and $i=1,2,\ldots,n$. For instance, if we focus on the value of function $u_{i}$ at point $x_{0}$, then $f(u_{i})=u_{i}(x_{0})$. ESS is calculated by the following formula:
$\text{ESS}=\text{ESS}(f,\{u_{k}\}_{k=1}^{n})=n\left[1+2\sum_{k\geq 0}\gamma_{k}(f)\right]^{-1},$
where $n$ is the number of samples and $\gamma_{k}(f)$ denotes the autocorrelation function of the samples, which is estimated by:
$\gamma_{k}(f)=\frac{c_{k}}{c_{0}},\quad c_{k}=\frac{1}{n}\sum_{s=\max(1,-k)}^{\min(n-k,n)}[f(u_{s+k})-\overline{f}][f(u_{s})-\overline{f}],$
where $\overline{f}=\frac{1}{n}\sum_{s=1}^{n}f(u_{s})$. In the discussion of all simple smooth problems, when we discuss ESS, the corresponding functional $f$ is defined as $f(u)=u(0.5)$.
For inverse problem (4.3), we employed the pCN algorithm with a step size $\beta=0.01$, collecting $10^{6}$ samples. Analyzing the trace plot, we discarded the initial $10^{5}$ samples and computed the ESS of the remaining samples as
$\text{ESS}_{p}=76,$
where $\text{ESS}_{p}$ denotes the ESS of the samples generated by pCN algorithm. Despite collecting a large number of samples, the $\text{ESS}_{p}$ was found to be relatively small. The results from the pCN algorithm were used as a benchmark. To ensure adequate sampling for pCN, we collected $3\times 10^{6}$ samples, and the ESS of the samples can be computed as $\text{ESS}_{p}=239$.
For comparison, we generate samples with the NF-iVI algorithm. Following Algorithm 1, we use 24 layers to illustrate the functional Householder flow, and the model are trained for 5000 iterations $(K=5000)$, with each iteration utilizing 30 samples $(N=30)$. We employ a step decay learning rate schedule with an initial learning rate $\alpha_{0}=0.01$ and a multiplicative factor $\tau=0.8$ after every 500 iterations. We collected 1000 samples and calculated the ESS as
$\text{ESS}_{\text{NF}}=926.$
Remarkably, even with only 1000 samples, the ESS of the samples proposed from NF-iVI still significantly exceeded that of pCN. Furthermore, additional samples can be easily generated without incurring substantial computational overhead. For example, once the training is finished, we can collect 10,000 samples using NF-iVI and compute the ESS as $\text{ESS}_{\text{NF}}=9658$.
@@
Firstly, we provide some discussions of the mean function of approximate posterior. In subfigure (a) of Figure 1, we compare the mean of the approximate posterior obtained by the NF-iVI method with the ground truth. The blue solid line represents the mean of NF-iVI estimate, while the red dashed line represents the ground truth. The shaded blue region denotes the $95\%$ credibility interval of the approximate posterior. We observe that the $95\%$ credibility interval encompasses the ground truth, reflecting the inherent uncertainty in the parameter estimation.
Subfigure (b) of Figure 1 presents a similar comparison between the mean of the estimated posterior measure obtained by the pCN algorithm and the ground truth. The blue shaded area again represents the $95\%$ credibility interval of the estimated posterior. As illustrated in the comparison between (a) and (b), the mean of the approximate posterior generated by functional Householder flow closely resembles the mean of posterior obtained from the pCN algorithm.
@@
To further support the effectiveness of estimating the mean of NF-iVI method, we provide numerical evidence. By comparing it with the pCN sampling algorithm, we calculate the relative error between their estimates:
[…]
@@
where $u_{N}^{*}$ denotes the estimated mean from functional Householder flow and $u_{p}^{*}$ represents the estimated mean from the pCN algorithm. Obviously, the small relative error suggests the effectiveness of the proposed algorithm.
@@
As a result, combining visual (Figure 1) and quantitative (relative error given in (4.5)) evidences, the NF-iVI method provides a good approximation of the mean of posterior.
(a) Result of Functional Householder flow
(b) Result of pCN algorithm
@@
Figure 1: A comparison of the approximate posterior obtained by functional Householder flow with 24 layers and the posterior generated by the pCN algorithm. The blue shade area represents the $95\%$ credibility region of estimated posterior. (a): Result of the approximate posterior generated by functional Householder flow compared with the background truth of $u$. (b): Result of the posterior generated by pCN algorithm compared with the background truth of $u$.
@@
Secondly, we offer a discussion on the approximation of the posterior covariance function. Our emphasis is on presenting evidence that demonstrates the NF-iVI method’s capability to provide a robust estimate of the covariance. To facilitate these comparisons, we need to introduce the definition of variance function and covariance function.

### v2

@@ 3 Conditional Normalizing Flow in Function Space
While Section 2 establishes the theoretical framework of functional normalizing flow, the practical application of Algorithm 1 is hindered by the computational cost of repeatedly solving forward problems. To address this limitation, strategies for reducing the algorithm’s iteration are essential. In order to mitigate the computational burden, we propose conditional functional normalizing flow, which involve pre-training a model to map measurement data and their corresponding measurement points to the parameters of a functional normalizing flow, enabling efficient estimation of posterior for new measurement (similar techniques have also been employed for normalizing flows in Euclidean space ardizzone2019guided, winkler2019learning).
Our objective is to learn the map from measurement information (measurement data $\bm{d}$ and their corresponding measurement points $\bm{x}$) to the parameters of a functional normalizing flow, enabling approximation of posterior corresponding to the given measurement information. Following model training, the conditional functional normalizing flow can efficiently generate approximate posterior for new measurement information. While this provides an initial estimate, further refinement can be achieved through fine-tuning the model on specific measurement information, leveraging the initial estimate as a starting point. This approach significantly reduces the number of iterations required compared with training from a random initialization.
When constructing conditional functional normalizing flow, the position of the measurement points $\bm{x}$ and the measurement data $\bm{d}$ both play important roles. Therefore, we need to reformulate the framework proposed in Section 2.
Let $\bm{x}=(x^{1},x^{2},\ldots,x^{N_{\bm{d}}})$ represents the measurement points of the PDE solution $w$, $\bm{d}$ denotes the measurement data, and let $\mathcal{G}\mathrel{\mathop{\ordinarycolon}}\mathcal{H}_{u}\rightarrow\mathcal{H}_{w}$ represents the PDE solution operator, where $\mathcal{H}_{u}$ is the space of parameter, and $\mathcal{H}_{w}$ is the space of solution. The measurement points, denoted by $\bm{x}$, influence the measurement operator $\mathcal{S}_{\bm{x}}$:
$\mathcal{S}_{\bm{x}}\mathrel{\mathop{\ordinarycolon}}\mathcal{H}_{w}\rightarrow\mathbb{R}^{N_{\bm{d}}},$
@@
The inverse problem can be formulated as
$\bm{d}=\mathcal{S}_{\bm{x}}\mathcal{G}(u)+\bm{\epsilon},$
where $\bm{d}\in\mathbb{R}^{N_{\bm{d}}}$ is the measurement data, $u\in\mathcal{H}_{u}$ is the parameter of interest, and $\bm{\epsilon}$ is a Gaussian random vector with zero mean and covariance matrix $\bm{\Gamma}_{\text{noise}}\mathrel{\mathop{\ordinarycolon}}=\tau^{-1}\textbf{I}$ ($\tau$ is a fixed positive number, I denotes the $N_{\bm{d}}$-dimensional identity matrix), which means
@@
Our objective is to infer the possible values of the model parameter $u$ based on the measurement data $\bm{d}$ and their corresponding measurement points $\bm{x}$ by the corresponding Bayes’ formula:
@@
where $\mu_{(\bm{x},\bm{d})}$ is the posterior, and the potential function $\Phi_{(\bm{x},\bm{d})}\mathrel{\mathop{\ordinarycolon}}\mathcal{H}_{u}\rightarrow\mathbb{R}$ is defined as
$\Phi_{(\bm{x},\bm{d})}(u)=\frac{1}{2}\lVert\bm{d}-\mathcal{S}_{\bm{x}}\mathcal{G}(u)\rVert^{2}_{\bm{\Gamma}_{\text{noise}}}$
with $\lVert\cdot\rVert_{\bm{\Gamma}_{\text{noise}}}\mathrel{\mathop{\ordinarycolon}}=\lVert\bm{\Gamma}_{\text{noise}}^{-1/2}\cdot\rVert$, and $Z_{\mu}^{(\bm{x},\bm{d})}$ is the normalization constant given by
@@
The goal of conditional functional normalizing flows is to approximate the posterior distribution defined by Bayes’ formula, where the forward operator $\mathcal{G}$ is fixed while the measurement information $(\bm{x},\bm{d})$ varies. By entering the measurement pair $(\bm{x},\bm{d})$ into the neural network, we aim to directly construct an approximate probability measure $\mu_{(\bm{x},\bm{d})}$ for the corresponding posterior distribution, as defined by Bayes’ formula (3.1). It is important to note that the measurement information $(\bm{x},\bm{d})$ is both dynamic and heterogeneous. Variations in the number, positions, and values of measurement points lead to varying dimensions and scales of the information. To facilitate the application of neural networks, we therefore seek to transform this heterogeneous information into a fixed-length vector representation.
@@
$\mathcal{S}_{\bm{x}}^{*}\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{N_{\bm{d}}}\rightarrow\mathcal{H}_{w}^{*}.$
@@
Unlike functional normalizing flow introduced in Section 2, the conditional network $\mathcal{N}_{\lambda}(\bm{v})$ propose a novel approach where the model parameters for each measurement information $(\bm{x},\bm{d})$ are generated directly, rather than being learned through training. In other words, once the parameters $\lambda$ of the neural network $\mathcal{N}_{\lambda}(\bm{v})$ are trained, we can directly generate the parameters of the corresponding flow model for any given measurement information $(\bm{x},\bm{d})$. Next, we detail the training process for the parameters $\lambda$ of the network $\mathcal{N}_{\lambda}(\bm{v})$.
The network $\mathcal{N}_{\lambda}(\bm{v})$ is trained on the dataset $D_{train}=\{(\bm{x}_{i},\bm{d}_{i})\}_{i=1}^{N_{train}}$, where $({\bm{x}}_{i},{\bm{d}}_{i})$ represents a pair of measurement data $\bm{d}_{i}$ and its corresponding measurement points $\bm{x}_{i}$ for $i=1,2,\ldots,N_{train}$. For a specific measurement information pair $(\bm{x}^{*},\bm{d}^{*})$, the parameters of the corresponding functional normalizing flow are given by $\theta^{*}=\mathcal{N}_{\lambda}(\bm{v}^{*})$, where $\bm{v}^{*}=\mathcal{W}(\bm{x}^{*},\bm{d}^{*})$. As shown in (3.1), we denote the posterior corresponding to $(\bm{x}^{*},\bm{d}^{*})$ as $\mu_{(\bm{x}^{*},\bm{d}^{*})}$. We hypothesize that the flow model $f_{\theta^{*}}$ can transform the prior $\mu_{0}$ into a measure $\nu_{\lambda}(\bm{v}^{*})$ that closely approximates the posterior $\mu_{(\bm{x}^{*},\bm{d}^{*})}$. Consequently, we expect the Kullback-Leibler divergence $D_{\text{KL}}(\nu_{\lambda}(\bm{v}^{*})||\mu_{(\bm{x}^{*},\bm{d}^{*})})$ to be small. Let $\{\bm{v}_{i}=\mathcal{W}(\bm{x}_{i},\bm{d}_{i})\}_{i=1}^{N_{train}}$ denote the vectors corresponding to the dataset $D_{train}$, the overall error function can then be written as
@@
For a detailed description of the algorithm, we refer the reader to Algorithm 2.
@@
1: Initialize the neural network parameter $\lambda$ to $\lambda_{0}$, the learning rate schedule to $\alpha_{k}$, the training steps to $K$, the sampling size to $M$, and the training dataset to $D_{train}=\{(\bm{x}_{1},\bm{d}_{1}),(\bm{x}_{2},\bm{d}_{2}),\ldots,(\bm{x}_{N_{train}},\bm{d}_{N_{train}})\}$. Then, initialize the iteration counter $k$ to 0;
@@
The neural network $\mathcal{N}_{\lambda}(\bm{v})$ serves as a computationally efficient estimator for the initial posterior approximation. For a specific measurement pair $(\bm{x}^{*},\bm{d}^{*})$, the input vector $\bm{v}^{*}=\mathcal{W}(\bm{x}^{*},\bm{d}^{*})$ produces the initial flow parameters $\theta^{*}=\mathcal{N}_{\lambda}(\bm{v}^{*})$. The corresponding normalizing flow $f_{\theta^{*}}$ approximates the true posterior $\mu_{(\bm{x}^{*},\bm{d}^{*})}$ by pushing forward the prior to the measure $\nu_{\lambda}(\bm{v}^{*})$. However, due to the inherent capacity limits of $\mathcal{N}_{\lambda}(\bm{v})$, this initial estimate may deviate from the target posterior, potentially leading to inaccurate uncertainty quantification. To refine this approximation, we propose a sample-specific fine-tuning strategy: the inverse problem associated with $(\bm{x}^{*},\bm{d}^{*})$ is further optimized via Algorithm 1 with the initialization $\theta_{0}=\theta^{*}$. As demonstrated in Section 5, this informed initialization significantly accelerates convergence compared to random starts.
@@ 4 Numerical Examples of Functional Normalizing Flow
To demonstrate the effectiveness of our novel framework for solving inverse problems, we present three illustrative examples, i.e., the simple elliptic equation, the steady-state Darcy flow equation and the electrical impedance tomography (EIT) problem. These examples enable a straightforward visualization of the approximate posterior distribution. The code used for simulations is available on GitHub at https://github.com/jjx323/FunctionalNormalizingFlow. All programs ran on a system with a 13th Gen Intel(R) Core(TM) i7-13700, NVIDIA GeForce RTX 4090, and Ubuntu 20.04.5 LTS.
In the first example, given the linear nature of the inverse problem, we utilize functional Householder flow and functional projected transformation flow as our flow models. For the second and third examples, which involves a nonlinear inverse problem, we employ functional planar flow and functional Sylvester flow. These flow models are specifically designed to capture intricate nonlinear relationships, enabling the transformation of a Gaussian distribution into a non-Gaussian distribution that more closely approximates the posterior distribution associated with the nonlinear inverse problem.
@@ 4.1 Simple Elliptic Equation
$\begin{split}-\beta\Delta w+w&=u,\quad\text{in}\ \Omega,\\ \frac{\partial w}{\partial\bm{n}}&=0,\quad\text{on}\ \partial\Omega,\end{split}$ (4.1)
where $\Omega=(0,1)\subset\mathbb{R}$, $\beta>0$ is a positive constant ($\beta=0.01$ in our experiment), and $\bm{n}$ denotes the outward unit normal vector. The forward operator is defined as
@@
    Let the domain $\Omega$ be an interval $(0,1)$ with $\partial\Omega=\{0,1\}$, and the measurement data are assumed to be $\{w(x^{i})|i=0,1,\ldots,10\}$, where $x^{i}=i/10$.
@@
    In order to avoid inverse crime kaipio2006statistical, the data is generated on a fine mesh with the number of grid points equal to $10^{4}$. And we use different sizes of mesh $n=\{50,75,100,200,300\}$ in the inverse stage.
Since the inverse problem under consideration is linear, this subsection demonstrates the efficacy of two linear flows—the functional Householder flow and the functional projected transformation flow—as well as the discretization invariance of the proposed model. We assume that the data produced from the underlying ground truth
@@
To demonstrate the effectiveness of our proposed algorithm, we conduct a comparative experiment with the preconditioned Crank-Nicolson (pCN) algorithm. It has been theoretically established that the pCN algorithm can generate samples from the true posterior distribution dashti2013bayesian. Consequently, we treat the samples obtained from pCN as benchmark results for comparison. However, as a Markov chain Monte Carlo (MCMC) method, pCN samples often exhibit high autocorrelation. To acquire a sufficient number of independent samples, numerous iterations are typically required, and the computational cost of pCN scales directly with the desired number of independent samples. In contrast, the computational burden of NF-iVI primarily lies in the training process of the flow model. Once trained, the generated samples are mutually independent, which enables efficient sampling. When performing the pCN algorithm, we produce $3\times 10^{6}$ samples for each chain, and discard the first $10^{5}$ samples as burn-in when calculating quantities of interest (similar with pinski2015algorithms, cotter2013). It is notable that running the pCN algorithm required approximately 9 minutes, whereas training our FNF model took only 25 seconds.
@@
Next, we will compare the posterior distribution obtained by the two algorithms in detail to verify the effectiveness of functional normalizing flow. Firstly, we provide some discussions of the mean function of approximate posterior. In subfigure (a)(b) of Figure 1, we compare the mean of the approximate posterior obtained by the NF-iVI method with the ground truth. The blue solid line represents the mean of NF-iVI estimate, while the red dashed line represents the ground truth. The shaded green region denotes the $95\%$ credibility interval of the approximate posterior. We observe that the $95\%$ credibility interval encompasses the ground truth, reflecting the inherent uncertainty in the parameter estimation. Subfigure (c) of Figure 1 presents a similar comparison between the mean of the estimated posterior measure obtained by the pCN algorithm and the ground truth. As illustrated in the comparison between (a)(b) and (c), the mean of the approximate posterior generated by two linear flows closely resembles the mean of posterior obtained from the pCN algorithm.
@@
To further support the conclusion, we provide numerical evidence. By comparing NF-iVI with the pCN sampling algorithm, we calculate the relative error between their estimates:
$\textbf{Householder relative error:}=\frac{\lVert u_{HN}^{*}-u_{p}^{*}\rVert_{L^{2}}^{2}}{\lVert u_{p}^{*}\rVert_{L^{2}}^{2}}=0.00129,$ (4.5)
$\textbf{projected transformation relative error:}=\frac{\lVert u_{PN}^{*}-u_{p}^{*}\rVert_{L^{2}}^{2}}{\lVert u_{p}^{*}\rVert_{L^{2}}^{2}}=0.00271,$ (4.6)
@@
where $u_{HN}^{*}$ denotes the estimated mean of functional Householder flow, $u_{PN}^{*}$ denotes the estimated mean of functional projected transformation flow, and $u_{p}^{*}$ represents the estimated mean of pCN algorithm. Obviously, the small relative errors suggest the effectiveness of our proposed algorithm.
@@
As a result, combining visual (Figure 1) and quantitative (relative errors given in Equations (4.5), (4.6)) evidences, the NF-iVI method provides a good approximation of the mean of posterior.
(a) Householder flow
(b) projected transformation flow
(c) pCN algorithm
@@
Figure 1: A comparison of the approximate posterior obtained by functional Householder flow with 24 layers Householder flow, 5 layers functional projected transformation flow and the posterior generated by the pCN algorithm. The green shade area represents the $95\%$ credibility region of estimated posterior. (a)(b): Result of the approximate posterior generated by functional Householder flow, functional projected transformation flow compared with the background truth of $u$. (c): Result of the posterior generated by pCN algorithm compared with the background truth of $u$.
@@
Secondly, we offer a discussion on the approximation of the posterior covariance function. The emphasis is placed on demonstrating that the NF-iVI method provides robust covariance estimates. To facilitate these comparisons, we need to introduce the definition of the variance and the covariance function.